One-eighth of Jesse Blacks inventory was destroyed by fire. He sold the remaining part, which was slightly damaged, for four-fifths of its value and received $3800. (a) What was the value of the destroyed part of the inventory? (b) What was the value of the inventory before the fire? (a) The value is s Round to the nearest cent as needed.) (b) The value is s Round to the nearest cent as needed.)
Answer #1

1. Let the value of inventory be x.
Inventory partially damaged = 1-(1/8) = 7/8 th of x
Partially damaged inventory sold for 4/5 of price which amounts to $3800
Therefore, four-fifth of seven-eights of x = $3800
or, (4/5)*(7/8)*x = $3800

7x/10 = 3800
x = 3800*10/7 = $ 5429

(a) value of inventory destroyed = 1/8th of $5429 = (1/8)*5429= $ 679
(b) value of inventory before fire = x = $5429

Thanks

Let me know if it's wrong